OpenStudy (anonymous):

Solve for d. sqrt--d^2+1 =1-d

OpenStudy (angela210793):

\[\sqrt{d^2-1}=1-d\]

OpenStudy (anonymous):

\[is this: \sqrt{d ^{2}+1}=1-d ?\]

OpenStudy (anonymous):

yes angela thats the exact problem (:

OpenStudy (anonymous):

if it is : \[d ^{2}+1=(1-d)^{2}\] solve it for d

OpenStudy (anonymous):

angela has the correct format of the problem

OpenStudy (anonymous):

\[d ^{2}+1=1-2d+d ^{2}\]

OpenStudy (anonymous):

d=1

OpenStudy (angela210793):

d^2-1=(1-d)^2 d^2-1=1-2d-d^2 d^2-1-1+2d+d^2=0 2d^2+2d-2=0 d^2+d-1=0 D=b^2-4ac=1-4*1*(-1)=5 d1=(-1-sqr5)/2 d2=(sqr5-1)2

OpenStudy (angela210793):

1d2=(-b+-sqrD)/2a

OpenStudy (anonymous):

\[d ^{2} - canceled out as well as 1\] d=0

OpenStudy (anonymous):

inik is right

OpenStudy (anonymous):

angela21073.. per original: should be d^2 +1 (not -)...?

OpenStudy (anonymous):

it is 0

OpenStudy (anonymous):

good :)

OpenStudy (anonymous):

d=0

OpenStudy (angela210793):

sorry didn't get it...where? can u write it?

OpenStudy (angela210793):

I got it

OpenStudy (angela210793):

:(:(:( All that for nothing :(:(:(:( Sorry Jazmine :(

OpenStudy (anonymous):

its okay angela u have helped me alot before too (: its just one mistake